OpenStudy (anonymous):

find the area inside the cardioid r=1-sin theta and outside the circle r=1/2

OpenStudy (anonymous):

you just want an answer?

OpenStudy (anonymous):

I would definitely exploit the symmetry so you only have one definite integral

OpenStudy (anonymous):

i wanted to c how to do it

OpenStudy (anonymous):

the integral you want is\[\int^{\frac{3\pi}{2}}_{\frac{5\pi}{6}}\left[\left[1-\sin\theta\right]^2-1^2\right]d\theta\]

OpenStudy (anonymous):

i am exploiting the symmetry

OpenStudy (anonymous):

it is best to sketch the region for these problems

OpenStudy (anonymous):

sorry that 1 should be 1/2

OpenStudy (anonymous):

and one of your limits should be pi/6

OpenStudy (anonymous):

well

OpenStudy (anonymous):

what is the lower limit then?

OpenStudy (anonymous):

its pi/6

OpenStudy (anonymous):

what is you upper limit then?

OpenStudy (anonymous):

5pi/6

OpenStudy (anonymous):

are you looking at the graph of the region?

OpenStudy (anonymous):

if you use those limits you will capture the wrong area

OpenStudy (anonymous):

hey i figured it out ty

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

those are tricky, the iterated integrals in multi-variable are even more awesome

OpenStudy (anonymous):

yah. oh do you think you can help me with graphing a problem u can see the problem if u click on this link http://www.twiddla.com/511680